OpenStudy (anonymous):

1. What is the factored form of 4x^2 + 12x + 5?

OpenStudy (aroub):

(2x+5)(2x+1)

OpenStudy (hoblos):

4x^2 + 12x + 5 = 4x^2 +2x+10x+5 = 2x(2x+1) + 5(2x+1) = (2x+1)(2x+5)

OpenStudy (anonymous):

(2x+5)(2x+1)